OpenStudy (anonymous):

Someone please help, even if just for one of them! 1.Write the equation of the line with a slope of -5 and a y-intercept of (0,3). 2.Write the equation of the line with a slope of -1/3 and passing through the point (6, -4). 3.Write the equation of the line passing through the points (0, -4) and (-2, 2). 4.Write the equation of the line passing through the points (-6,1) and (-4,2). 5.Write the equation of the line with an undefined slope, passing through the point (2, 5).

jimthompson5910 (jim_thompson5910):

how far did you get?

OpenStudy (anonymous):

I don't know how to start

jimthompson5910 (jim_thompson5910):

let's start with #1

jimthompson5910 (jim_thompson5910):

1.Write the equation of the line with a slope of -5 and a y-intercept of (0,3).

jimthompson5910 (jim_thompson5910):

The given slope is -5, so m = -5 (m is the slope) the y-intercept is (0,3) or just 3, so b = 3

jimthompson5910 (jim_thompson5910):

y = mx + b ... start with slope intercept form y = -5x + b ... replace m with -5 (since m = -5) y = -5x + 3 ... replace b with 3 (since b = 3)

jimthompson5910 (jim_thompson5910):

so the equation with a slope of -5 and a y-intercept of (0,3) is y = -5x+3

jimthompson5910 (jim_thompson5910):

make sense?

OpenStudy (anonymous):

Okay thank you! so do you use the same formula for all of them?

jimthompson5910 (jim_thompson5910):

yes, but the ones that don't have a given slope, you need to find it first

jimthompson5910 (jim_thompson5910):

and if you don't know the intercept, you need to find that as well

jimthompson5910 (jim_thompson5910):

once you know the slope and y-intercept, you know the equation based on what is done above

OpenStudy (anonymous):

okay so what if they give me a line on a graph and tell me to write the equation?

jimthompson5910 (jim_thompson5910):

identify two points on that line find the slope using those two points then use one of those points (along with the slope) to find the y-intercept

jimthompson5910 (jim_thompson5910):

to make life easier, you can visually look for the y-intercept so you don't have to do much work

jimthompson5910 (jim_thompson5910):

does that make sense?

OpenStudy (anonymous):

kind of

jimthompson5910 (jim_thompson5910):

what are two points on that line?

OpenStudy (anonymous):

-1 and 3

jimthompson5910 (jim_thompson5910):

write them as ordered pairs for instance, the point (3,1) is on the line (see attached image)

jimthompson5910 (jim_thompson5910):

what other point is on the line?

OpenStudy (anonymous):

(-3,-3)

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

now use these 2 points to find the slope

jimthompson5910 (jim_thompson5910):

you do so by using the slope formula \[\large m = \frac{y_{2}-y_{1}}{x_{2}-x_{1}}\]

OpenStudy (anonymous):

okay!

jimthompson5910 (jim_thompson5910):

tell me what you get

OpenStudy (anonymous):

,m = -3-1/-3-3

jimthompson5910 (jim_thompson5910):

keep going

OpenStudy (anonymous):

-2

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

-3 - 1 is equal to ???

OpenStudy (anonymous):

-4 /-6

jimthompson5910 (jim_thompson5910):

reduce that

OpenStudy (anonymous):

-2/-3

jimthompson5910 (jim_thompson5910):

keep going

jimthompson5910 (jim_thompson5910):

two negatives divide to get a _____

OpenStudy (anonymous):

positive

jimthompson5910 (jim_thompson5910):

so the slope is 2/3

jimthompson5910 (jim_thompson5910):

so given any two points, you can find the slope of the line through those two points (using the slope formula) so you'll be using this formula for #3 and #4

jimthompson5910 (jim_thompson5910):

we now have a slope of 2/3, so m = 2/3

jimthompson5910 (jim_thompson5910):

how do we find the y-intercept? ie how do we find the value of b?

OpenStudy (anonymous):

i dont know

jimthompson5910 (jim_thompson5910):

remember we have the general equation y = mx+b

jimthompson5910 (jim_thompson5910):

we know the slope is m = 2/3 we also know that the point (-3,-3) is on the line so this means that x = -3 and y = -3

jimthompson5910 (jim_thompson5910):

plug all this in, then solve for b \[\large y = mx+b\] \[\large y = \frac{2}{3}x+b\] \[\large -3 = \frac{2}{3}(-3)+b\] \[\large -3 = \frac{2}{3}*\frac{-3}{1}+b\] \[\large -3 = \frac{2*(-3)}{3*1}+b\] \[\large -3 = \frac{-6}{3}+b\] \[\large -3 = -2+b\] \[\large b = ???\]

OpenStudy (anonymous):

-5?

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

you have to add 2 to both sides to isolate b

OpenStudy (anonymous):

oh crao i added them, it would be -1

jimthompson5910 (jim_thompson5910):

good, look at the graph and notice where it crosses the y axis

OpenStudy (anonymous):

oh wow thats alot easier hah

jimthompson5910 (jim_thompson5910):

yes BUT you won't always have the graph so it helps to know how to do this without a graph

jimthompson5910 (jim_thompson5910):

besides, the line may intersect the y-axis at some point that's not a whole number

jimthompson5910 (jim_thompson5910):

anyways, we know the slope is 2/3 the y-intercept is -1 so this means m = 2/3 b = -1

jimthompson5910 (jim_thompson5910):

Now plug these pieces of info in \[\large y = mx+b\] \[\large y = \frac{2}{3}x+b\] \[\large y = \frac{2}{3}x+(-1)\] \[\large y = \frac{2}{3}x-1\] So the equation of the line is \[\large y = \frac{2}{3}x-1\]

OpenStudy (anonymous):

thank you !

jimthompson5910 (jim_thompson5910):

you're welcome